A food product contains 1 x 106 CFU/g of Escherichia coli. Heating at 80 C for 10 minutes will reduce this population to 1 x 104 CFU/g. The z-value of this E. coli strain is 3 C. How long would this food product have to be heated at 89 C to achieve the same microbial destruction?
Answer #1

The population of E. Coli initially (a) = 1 x 106 CFU/g

The population of E. Coli after heating (b) = 1 x 104 CFU/g.

time = 10 minutes

temperature = 80 degrees Celcius.

D80 = time/ (log a - log b) = 10/ (log 106 - log 104) = 10/ (6 -4) = 10/2 = 5.

The D-value at 80 degrees Celcius is 5 minutes.

Z value = (T2 - T1)/ (log D2 - log D1)

3 = (89 - 80)/ (log D2 - 5)

log D2 - 5 = 9/3 = 3

log D2 = 3 + 5 = 8

hence, to achieve the same microbial destruction, the food would have to be heated at 89 degrees celcius for 8 minutes.